# SAR-Net: Multi-scale Direction-aware SAR Network via Global Information Fusion

### v2

@@ IV-B Accuracy Metrics
TABLE VI: Comparison with Other Methods on SAR-AIRcraft-1.0. The Accuracy of Each category is Represented by mAP with an Intersection Over the Union Threshold of 0.5
Method A220 A320/321 A330 ARJ21 Boeing737 Boeing787 other mAP50(%) F1-score Params(M) GFLOPs
Faster R-CNN[57] 77.3 97.0 86.4 74.0 59.5 74.0 71.2 77.2 0.746 41.09 142.73
Cascade R-CNN[58] 75.2 97.7 88.6 79.5 55.0 69.6 69.8 76.9 0.728 69.11 185.89
YOLOv6n[47] 92.6 97.5 96.2 81.4 78.3 80.6 78.9 86.5 0.813 4.64 17.80
YOLOv8n[59] 90.2 97.8 96.5 84.3 81.7 79.9 83.5 87.7 0.830 3.15 13.57
SKG-Net[60] 66.0 79.9 79.2 66.1 64.8 70.4 72.1 71.2 0.659 31.38 103.69
@@
MLSDNet[61] 85.1 96.9 91.5 83.2 71.7 72.1 78.4 82.7 0.779 1.42 6.72
HRLE-Det[62] 90.6 98.1 96.7 82.6 78.9 79.1 82.5 86.9 0.840 3.02 14.20
SAR-Net 91.0 98.5 96.9 92.6 86.6 82.4 83.6 90.2 0.856 6.09 20.80
@@
TABLE VII: Comparison with the Latest Object Detection Methods and SAR Ship Detection Methods on SSDD
Method mAP50(%) F1-score Params(M) GFLOPs
Faster R-CNN[57] 96.5 0.953 41.09 91.33
Cascade R-CNN[58] 97.0 0.955 69.11 118.90
YOLOv6n[47] 97.9 0.949 4.64 11.40
YOLOv8n[59] 98.1 0.956 3.15 8.70
LPEDet[63] 97.5 0.938 5.66 18.34
MLSDNet[61] 97.2 0.937 1.42 4.31
HRLE-SARDet[62] 97.6 0.948 3.02 9.10
YOLOv8m[59] 98.4 0.968 25.40 78.92
SAR-Net 98.8 0.972 6.09 13.76
IV-C Ablation Study
To comprehensively and detailedly analyze the effectiveness of SAR-Net, we conducted extensive ablation experiments on the SAR-AIRcraft-1.0 dataset with different framework designs and internal modules. The best results are highlighted in bold.
IV-C1 Impact of Unity Compensation Mechanism
To study the effectiveness of each module in the unity compensation mechanism and verify the correctness of uniting deep and shallow layers’ global information, we conducted ablation experiments on three of the modules and measured the performance with mAP50 and F1-score, a common indicator in object detection. In addition, Parameter sizes and GFLOPs are also shown in Table II. We use YOLOv6n [47] as the baseline for ablation experiments to embed the modules of our unity compensation mechanism one by one. The feature alignment process is included in all experiments except the baseline. It can be seen that only when the Deep-MFM module is used, the accuracy is lower than the baseline. This is because it pays more attention to large objects, while most of the SAR images are small and medium-sized objects. In addition, other experiments have achieved performance improvements without significantly increasing the number of calculations, which proves the effectiveness of our global information fusion idea. In particular, when the MEM module is not used, we use simple addition to compensate for the original features.
@@ IV-C1 Impact of Unity Compensation Mechanism
TABLE VIII: Comparison with the Latest Object Detection Methods and SAR Ship Detection Methods on HRSID
Method mAP50(%) F1-score Params(M) GFLOPs
Faster R-CNN[57] 81.0 0.742 41.09 142.73
Cascade R-CNN[58] 82.7 0.813 69.11 185.89
YOLOv6n[47] 91.2 0.873 4.64 17.80
YOLOv8n[59] 91.4 0.877 3.15 13.57
LPEDet[63] 89.9 0.862 5.66 28.61
MLSDNet[61] 89.1 0.845 1.42 6.72
HRLE-SARDet[62] 90.8 0.866 3.02 14.20
CMFT[64] 89.6 0.859 - -
YOLOv8m[59] 92.8 0.891 25.40 123.12
SAR-Net 93.4 0.898 6.09 20.80
IV-C2 Impact of Direction-aware Attention Module
To prove the effectiveness of the proposed direction-aware attention module, we conducted a series of detailed experiments, as shown in Table III. We split the attention modules in two directions and embedded them in the YOLOv6n backbone separately, and both received performance improvements. In addition, we tried to embed DAM in different positions of the backbone. Except for the embedding position shown in Fig. 3, we also tried to move the DAM behind the RELU and each ERBlock, and obtained 0.9% and 0.6% performance improvement respectively with a small increase in parameters. Exhaustive experiments allowed us to determine the most suitable embedding position. These experimental results show that the acquisition of direction-aware information is significantly helpful for SAR images containing directional objects.
IV-C3 Collaboration of the Framework
The effectiveness of the components that make up SAR-Net has been demonstrated, and to further explore the synergy between them, we combined them and found the desired results. For the unity compensation mechanism, the backbone that can extract direction-aware information gives full play to the advantages of global fusion and enriches the information of multi-scale features. For the DAM-embedded backbone, the unity compensation mechanism retains detailed information and enhances the efficiency of information fusion and transmission. As shown in Table IV, after combining the two, their collaborative complementarity is fully utilized and performance is further improved by 0.4%.
@@ IV-C3 Collaboration of the Framework
IV-C4 Comparison of Different FPN Architecture
We chose the classic FPN structure and its latest variants AFPN [43] and RepBi-PAN [47] for comparison. We conducted the experiments based on YOLOv6n and the results are shown in Table V. There is a large accuracy gap when the parameters are similar. As we discussed in the section I, sequential local fusion will lead to information loss, especially for SAR images with large object scale variants. Top-down and bottom-up structures will lose a lot of detailed information and semantic information before detection. Global information fusion and compensation will retain the multi-scale information of the object to the greatest extent, therefore unity compensation mechanism achieves the best performance.
IV-D Comparisons with Previous Methods
To demonstrate the effectiveness and robustness of our proposed method, in addition to comparing it with classic and latest object detection methods on three datasets (SAR-AIRcraft-1.0, SSDD, HRSID), we also selected methods specially designed for SAR images. With the exception of the unreleased and non-detailed network structure of the comparative method, we successfully reproduced all other comparative methods. All methods are compared and evaluated from four aspects: mean Average Precision, F1-score, parameter size, and GFLOPs. The experimental parameter settings for each dataset remain consistent.
@@ IV-D Comparisons with Previous Methods
IV-D1 Comparison Results on SAR-AIRcraft-1.0
We used the SAR-AIRcraft-1.0 dataset to train and test several common benchmark detection methods, including two-stage detectors: Faster R-CNN [57], Cascade R-CNN [58], and one-stage detectors: YOLOv6 [47], YOLOv8 [59]. In addition to this, we also selected three latest SAR aircraft detection methods: SKG-Net [60], SA-Net [53] and MLSDNet [61], and a formidable model, HRLE-SARDet[62]. Table VI shows the detection results under different methods.
In the Faster R-CNN method, the accuracy is 77.2%, which shows to a certain extent that the SAR-AIRcraft-1.0 dataset has a certain detection difficulty because aircraft objects of different categories have similar structures and sizes, and the same category also presents multi-scale characteristics. As shown in Table VI, it can be seen that due to the fusion of global information, SAR-Net achieves the highest performance in almost every category, and on this basis, we only add a small amount of computational overhead, but the performance has been greatly improved. Considering the advanced nature of YOLOv8, we also experimented with its variants from small to large versions. We observed that the performance fluctuated within a range of 0.3%, with a significant increase in computational cost. In addition, the detection accuracy of each category has certain differences. For example, A320/321 performs best in different algorithms compared to other categories. This is mainly due to the special size of A320/321, with a fuselage length of more than 40 meters, making it easy to distinguish. For some objects in other categories, such as ARJ21 and Boeing aircraft, due to their multi-scale characteristics and varied directions, this pose a challenge for the network to learn invariant features within the same category. We enhance the detection capabilities of multi-scale and directional information-rich objects by fusing global information and performing compensation in deep and shallow layers respectively, as well as learning object direction-aware information.
IV-D2 Comparison Results on SSDD
To verify the superiority of our method on different datasets, we selected the widely used SSDD dataset and compared it with some of the latest object detection methods and SAR ship detection methods. A variety of experiments ensured the rigor of our work.
As indicated in the data presented in Table VII, our method exhibits strong competitiveness even when transferred to ship detection tasks. It can be observed that due to the relatively small size of the SSDD dataset, the current methods have already achieved a high level of average precision. However, due to the continued presence of multi-scale objects and interference from background noise, this has prevented further performance improvements in other methods. Note that our method excels in information fusion and transmission, making it less prone to influence by these problems and less prone to overfit on small datasets. Consequently, this further elevates the benchmark performance of the dataset. Additionally, in comparison to the state-of-the-art (SOTA) model YOLOv8’s larger version, we still achieve a 0.4% precision advantage on mAP50, while reducing parameters and GFLOPS to one-fourth and one-sixth, respectively.
IV-D3 Comparison Results on HRSID
To further validate the generalization and robustness of our method, we opted for the more challenging ship detection dataset, HRSID, in comparison to SSDD. HRSID presents greater difficulty with a higher number of instances, larger image resolutions, and significant variations in object scales. This dataset serves as a more stringent test of the model’s capabilities in feature extraction and utilization.
Continuing our evaluation, we chose state-of-the-art object detection methods, including those specifically tailored for ship detection in SAR images. Compared to other approaches, our method maintains excellent performance with a relatively small computational cost. Consistent with the results from SSDD, when compared to YOLOv8m, our method exhibits a 0.6% precision advantage on mAP50 and significantly smaller parameter size and GFLOPs, being one-fourth and one-sixth of YOLOv8m, respectively.
IV-E Result Visualization
@@
Due to the SAR-AIRcraft-1.0 dataset’s nature of fine-grained object detection, it presents certain detection challenges, leading to notable variations in detection results among different methods. To provide a more intuitive representation of the performance of various methods on SAR-AIRcraft-1.0, we selected a few images and visualized the detection results using different methods, as shown in Fig. 8. In comparison with other methods, our approach consistently identifies objects. While occasional false positives may occur, the incorporation of sufficient directional awareness information and the utilization of global fusion to minimize information loss result in minimal instances of false negatives in our detection results. For instance, the green bounding box (representing A220) in (b), the brown bounding box (representing ARJ21) in (c), and the orange bounding box (representing Boeing787) in (d) were undetected by other methods, highlighting the superiority of our approach in these scenarios.
@@
For the SSDD and HRSID datasets, given their high baseline levels of average precision and the single-category nature of object detection, we provide visualization results exclusively for SAR-Net. The visualizations for SSDD and HRSID are presented in Fig. 9 and Fig. 10, respectively. SAR-Net demonstrates excellent detection performance in both inshore and offshore backgrounds.
V Conclusion
@@
As the focus on deep learning-based remote sensing tasks grows, SAR image object detection has become a crucial research direction. Establishing a complementary relationship between high-level semantic information and low-level detailed information without losing critical details is essential for solving the multi-scale object feature fusion problem. In response to this, we propose a framework capable of globally fusing multi-scale direction-aware information, named SAR-Net. This framework effectively addresses the challenges of information loss due to sequential local fusion and the difficulty in capturing detailed information under complex background interference, allowing multi-scale detection features to focus on global information. Experiments conducted on SAR image aircraft and ship datasets demonstrate the effectiveness and excellent generalization of our framework. We believe this global information fusion perspective will provide valuable insights for SAR image detection tasks.

### v4

@@ 4.3.4 Comparison of different FPN architecture
Table 7: Comparison with other methods on SAR-AIRcraft-1.0. The accuracy of each category is represented by mAP with an intersection over the union threshold of 0.5.
Method A220 A320/321 A330 ARJ21 Boeing737 Boeing787 other mAP50(%) F1-score(%) Params(M) GFLOPs
Faster R-CNN[29] 77.3 97.0 86.4 74.0 59.5 74.0 71.2 77.2 74.6 41.0 142.7
Cascade R-CNN[30] 75.2 97.7 88.6 79.5 55.0 69.6 69.8 76.9 72.8 69.1 185.8
YOLOv6n[23] 92.6 97.5 96.2 81.4 78.3 80.6 78.9 86.5 81.3 4.6 17.8
YOLOv8n[31] 90.2 97.8 96.5 84.3 81.7 79.9 83.5 87.7 83.0 3.1 13.5
SKG-Net[32] 66.0 79.9 79.2 66.1 64.8 70.4 72.1 71.2 65.9 31.3 103.6
@@
MLSDNet[33] 85.1 96.9 91.5 83.2 71.7 72.1 78.4 82.7 77.9 1.4 6.7
HRLE-SARDet[34] 90.6 98.1 96.7 82.6 78.9 79.1 82.5 86.9 84.0 3.0 14.2
FBR-Net[35] 88.1 95.7 94.4 82.2 79.6 77.8 81.4 85.6 80.1 32.4 141.3
AFSar[36] 89.2 96.8 95.5 83.3 80.7 78.9 82.5 86.7 81.7 3.6 14.9
RT-DETR[37] 90.1 97.7 96.3 84.3 81.5 79.9 83.4 87.6 82.6 31.9 190.2
CCDN-DETR[38] 89.7 97.3 96.0 83.8 81.2 79.4 83.0 87.2 82.4 23.6 163.7
SAR-Net 91.0 98.5 96.9 92.6 86.6 82.4 83.6 90.2 85.6 6.0 20.8
4.4 Comparisons with previous methods
To showcase the effectiveness and robustness of our proposed method, we compared it with both classic and recent object detection methods on three datasets (SAR-AIRcraft-1.0, SSDD, HRSID). Additionally, we included comparisons with methods specifically designed for SAR images, such as those based on CNN and Transformer architectures. With the exception of the unreleased and non-detailed network structure of the comparative method, we successfully reproduce all other comparative methods. All methods are compared and evaluated from four aspects: mean Average Precision, F1-score, parameter size, and GFLOPs. The experimental parameter settings for each dataset remain consistent.
4.4.1 Comparison results on SAR-AIRcraft-1.0
We use the SAR-AIRcraft-1.0 dataset to train and test several common benchmark detection methods, including two-stage detectors: Faster R-CNN [29], Cascade R-CNN [30], and one-stage detectors: YOLOv6 [23], YOLOv8 [31]. In addition to this, we also selected three latest SAR aircraft detection methods: SKG-Net [32], SA-Net [25] and MLSDNet [33], and some formidable model, HRLE-SARDet[34], RT-DETR [37], CCDN-DETR [38], FBR-Net [35], AFSar [36]. Most of the methods are also applied to SSDD and HRSID. Table 7 shows the detection results under different methods.
@@ 4.4 Comparisons with previous methods
Table 8: Comparison with the latest object detection methods and SAR ship detection methods on SSDD.
Method mAP50(%) F1-score(%) Params(M) GFLOPs
Faster R-CNN[29] 96.5 95.3 41.0 91.3
Cascade R-CNN[30] 97.0 95.5 69.1 118.9
YOLOv6n[23] 97.9 94.9 4.6 11.4
YOLOv8n[31] 98.1 95.6 3.1 8.7
LPEDet[39] 97.5 93.8 5.6 18.3
MLSDNet[33] 97.2 93.7 1.4 4.3
HRLE-SARDet[34] 97.6 94.8 3.0 9.1
YOLOv8m[31] 98.4 96.8 25.4 78.9
FBR-Net [35] 94.1 92.8 32.4 94.0
AFSar [36] 97.1 94.7 3.6 9.9
RT-DETR [37] 97.5 95.7 31.9 125.3
CCDN-DETR [38] 97.2 95.6 23.6 109.1
SAR-Net 98.8 97.2 6.0 13.7
In the Faster R-CNN method, the accuracy is 77.2%, which shows to a certain extent that the SAR-AIRcraft-1.0 dataset has a certain detection difficulty because aircraft objects of different categories have similar structures and sizes, and the same category also presents multi-scale characteristics. As illustrated in Table 7, SAR-Net achieves the highest performance in nearly every category, primarily due to its integration of global information, and on this basis, we only add a small amount of computational overhead, but the performance has been greatly improved. Considering the advanced nature of YOLOv8, we also experimented with its variants from small to large versions. We observed that the performance fluctuated within a range of 0.3%, accompanied by a notable increase in computational cost. In addition, the detection accuracy of each category has certain differences. For example, A320/321 performs best in different algorithms compared to other categories. This is mainly due to the special size of A320/321, with a fuselage length of more than 40 meters, making it easy to distinguish. For some objects in other categories, such as ARJ21 and Boeing aircraft, due to their multi-scale characteristics and varied directions, this presents a challenge for the network in learning invariant features within the same category. We enhance the detection capabilities of multi-scale and directional information-rich objects by fusing global information and performing compensation in deep and shallow layers respectively, as well as learning object direction-aware information.
4.4.2 Comparison results on SSDD
To confirm the superiority of our approach on different datasets, we selected the widely used SSDD dataset and compared it against several state-of-the-art object detection methods, some of which were specially designed for SAR imagery. A variety of experiments ensured the rigor of our work.
@@ 4.4.2 Comparison results on SSDD
Table 9: Comparison with the latest object detection methods and SAR ship detection methods on HRSID.
Method mAP50(%) F1-score(%) Params(M) GFLOPs
Faster R-CNN[29] 81.0 74.2 41.0 142.7
Cascade R-CNN[30] 82.7 81.3 69.1 185.8
YOLOv6n[23] 91.2 87.3 4.6 17.8
YOLOv8n[31] 91.4 87.7 3.1 13.5
LPEDet[39] 89.9 86.2 5.6 28.6
MLSDNet[33] 89.1 84.5 1.4 6.7
HRLE-SARDet[34] 90.8 86.6 3.0 14.2
CMFT[40] 89.6 85.9 - -
YOLOv8m[31] 92.8 89.1 25.4 123.1
FBR-Net [35] 89.6 85.1 32.4 141.3
AFSar [36] 91.3 87.2 3.6 14.9
RT-DETR [37] 91.9 88.1 31.9 190.2
CCDN-DETR [38] 91.6 87.9 23.6 163.7
SAR-Net 93.4 89.8 6.0 20.8
As indicated in the data presented in Table 8, our method exhibits strong competitiveness even when transferred to ship tasks. It can be observed that due to the relatively small size of the SSDD dataset, the current methods have already achieved a high level of average precision. However, due to the continued presence of multi-scale objects and interference from background noise, this has prevented further performance improvements in other methods. Note that our method excels in information fusion and transmission, making it less prone to influence by these problems and less prone to overfit on small datasets. Consequently, this further elevates the benchmark performance of the dataset. Additionally, in comparison to the state-of-the-art (SOTA) model YOLOv8’s larger version, we still achieve a 0.4% precision advantage on mAP50, while reducing parameters and GFLOPS to one-fourth and one-sixth, respectively.
@@
4.4.3 Comparison results on HRSID
For additional validation of the generalization and robustness of our approach, we opted for the more challenging ship dataset, HRSID, in comparison to SSDD. HRSID presents greater difficulty with a higher number of instances, larger image resolutions, and significant variations in object scales. This dataset serves as a more stringent test of the model’s capabilities in feature extraction and utilization.
Table 10: Comparison with the latest oriented SAR ship object detection methods on SSDD and HRSID.
Method SSDD HRSID
mAP50(%) F1-score(%) mAP50(%) F1-score(%)
Yue[41] 90.7 88.3 86.5 83.9
Li[42] 96.5 94.8 86.3 83.2
Zhou[43] 93.5 91.6 87.8 84.7
KeyShip[44] 90.7 89.1 83.7 82.2
LPST-Det[45] 93.9 91.2 87.8 85.4
SAR-Net 98.8 97.2 93.4 89.8
Continuing our evaluation, we chose state-of-the-art object detection methods, some of which were specially designed for SAR imagery. Compared with other approaches, our method maintains excellent performance with a relatively small computational cost, as shown in Table 9. Consistent with the results from SSDD, when compared to YOLOv8m, our method exhibits a 0.6% precision advantage on mAP50 and significantly smaller parameter size and GFLOPs, being one-fourth and one-sixth of YOLOv8m, respectively.
@@ 4.4.3 Comparison results on HRSID
4.4.4 Comparison results on SSDD and HRSID with oriented bounding box detection methods
To further explore the direction-aware attention module for detecting rotated objects, we compared it with state-of-the-art SAR image rotation detection methods and conducted experiments on two ship datasets with rotation annotations. The results, as illustrated in Table 10, demonstrate that our method achieved the highest accuracy on both datasets. Despite not being specifically designed for rotation detection, the direction-aware attention module is still able to learn object feature information from two directions, thus assigning greater attention weights to the detected objects. Additionally, it retains this directional feature to provide perceptual information beneficial for detection.
4.5 Result visualization
@@
Due to the SAR-AIRcraft-1.0 dataset’s nature of fine-grained object detection, it presents certain detection challenges, leading to notable variations in detection results among different methods. To provide a more intuitive representation of the performance of various methods on SAR-AIRcraft-1.0, we selected a few images and visualized the detection results using different methods, as illustrated in Fig. 9. In comparison with other methods, our approach consistently identifies objects. While occasional false positives may occur, the incorporation of sufficient directional awareness information and the utilization of global fusion to minimize information loss result in minimal instances of false negatives in our detection results. For instance, the green bounding box (representing A220) in (b), the brown bounding box (representing ARJ21) in (c), and the orange bounding box (representing Boeing787) in (d) were undetected by other methods, highlighting the superiority of our approach in these scenarios.
@@
For the SSDD and HRSID datasets, given their high baseline levels of average precision and the single-category nature of object detection, we provide visualization results exclusively for SAR-Net. The visualizations for SSDD and HRSID are depicted in Fig. 10 and Fig. 11. SAR-Net demonstrates excellent detection performance in different backgrounds.
5 Conclusion
@@
As the focus on deep learning-based remote sensing tasks grows, SAR image object detection has become a crucial research direction. Establishing a complementary relationship between high-level semantic information and low-level detailed information without losing critical details is essential for solving the multi-scale object feature fusion problem. In response to this, we propose a framework capable of globally fusing multi-scale direction-aware information, named SAR-Net. This framework effectively addresses the challenges of information loss due to sequential local fusion and the difficulty in capturing detailed information under complex background interference, allowing multi-scale detection features to focus on global information. Experiments executed on SAR datasets of aircraft and ships demonstrate the effectiveness and excellent generalization of our framework. We believe this global information fusion perspective will provide valuable insights for SAR image detection tasks.
Declaration of Competing Interest
No conflict of interest exists in the submission of this manuscript, and all authors approve the manuscript for publication. We would like to declare that the work described was original research that has not been published previously, and is not under consideration for publication elsewhere.
Acknowledgments
This work was supported in part by the National Natural Science Foundation of China under Grant 62121001, Grant 62322117, Grant 62371365, and Grant U22B2014; and in part by the Young Elite Scientist Sponsorship Program through the China Association for Science and Technology under Grant 2020QNRC001.